OpenStudy (anonymous):

need some help please. I need to find the slope. 4y=x+8

OpenStudy (anonymous):

don't you just divide x+8 by 4 to get 1/4x+8/4?

OpenStudy (anonymous):

I am not sure that is why I am asking. I am stumped

OpenStudy (yrivers36):

4Y/4=X/4+8/4

OpenStudy (anonymous):

i see where I messed up

OpenStudy (yrivers36):

Y=1/4X+2 I BELIEVE

OpenStudy (anonymous):

i forgot the x/4 in my equation

OpenStudy (anonymous):

it is asking for an integer or a simplified fraction or is it undefined

OpenStudy (anonymous):

x/4 = 1/4 * x

OpenStudy (anonymous):

it is not looking for an equation. would it be 1/4

OpenStudy (anonymous):

or would it be 1/4x?

OpenStudy (anonymous):

1/4x because 1/4 is your x-slope

OpenStudy (anonymous):

ok thank you very much

OpenStudy (anonymous):

no problem :0]